OpenStudy (anonymous):

A cheetah runs along the path such that the displacement vs. time (meters vs. seconds) graph is given by the equation: x(t)=(6/π)sin-1t•cos-1t. What is the instantaneous velocity of the cheetah at the point (1/2, π/3)?

OpenStudy (dumbcow):

use derivatives of arcsin and arccos \[\frac{d}{dt}\sin^{-1} t = \frac{1}{\sqrt{1-t^{2}}}\] \[\frac{d}{dt}\cos^{-1} t = -\frac{1}{\sqrt{1-t^{2}}}\] Find v(t) = x'(t) , use product rule

OpenStudy (anonymous):

Thanks!